OpenStudy (anonymous):

The formula for an amount in an account with simple interest is A = P(1 + rt). Find r, if A = 384, P = 300, and t = 4. Answer r = 0.07 r = 0.035 r = 0.7 r = 0.35

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

384/3000=1+4r or 4r=384/300 -1 or r=84/(4*300) or r=84/1200=0.07